Astronomy Report – December 11th

Astronomy Report

Name:   Thomas Horn    Crew: 184
Date: 12/11/17

Sky Conditions: Clear

Wind Conditions: Still

Observation Start Time: 10:30am

Observation End Time: 1:30pm

Summary: It was reported there were interesting prominence’s at the limb of the sun.  These were not able to be observed, but two sunspots were identified in the below image. 

Objects Viewed: Sun, Sunspot

Problems Encountered: None

Operations Report – December 11th

Crew 184 Operations Report  12/11/2017

SOL: 10

Name of person filing report: HuntJ

Non-nominal systems:   EVA suits #5 / #6 (see below)

Notes on non-nominal systems: 

Generator (hours run): Generator turned off at 10:05 AM, nominal response.  Turned On at 5:45pm.

Solar— SOC @ 8:00am: 100%

              SOC @ 10:05am: 100%

             SOC @ 1:30 pm: 100%

             SOC @ 5:45pm:  65% (Generator Turned On)

Diesel –  99%

Propane –  Assume 69 percent volume (Due to trends from previous days)

Ethanol Free Gasoline (5 Gallon containers for ATV) –   1.5 Gallons

Water (trailer) –   400. gallons

Water (static) –   400 gallons

Trailer to Static Pump used –  Yes

Water (loft) – Static to Loft Pump used – Yes

Water Meter:  47 gallons

Toilet tank emptied: No

ATV’s Used: None,

Oil Added: No

ATV Fuel Used:  None

# Hours the ATVs were Used today: 00:00 Hours

Notes on ATVs: No ATVs used today due to no EVA.

Deimos rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

Sojourner rover used:  ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

HabCar used and why, where?  NO

General notes and comments: Power is back and we’re back in sim!!

Summary of internet: 1.0 GB of Bandwidth remaining

Summary of suits and radios:

Radios: Nominal after getting charged post HAB powerdown.  Crew has primarily retired old model radios to preferentially use the new radios delivered by Mission Control.  Communication is much crisper / easier to understand on new radios, particularly when on EVA.

EVA Suits:

EVA 5: Fan does not turn on.

EVA 6: Fan does not turn on.  One of the two black straps holding the Helmet Collar in place to the suit is missing, resulting in helmet sitting incorrectly when in use and forcing the EVA crewmembers head down.

EVA 8: Right latch securing the back panel is loose. Was able to re secure by threading a  zip tie through the latch and tieing it down. After a quick vibration test (read: Shaking the suit a little) everything held steady. EVA 8 is operational at this time.

Summary of Hab operations:  Nominal

Summary of GreenHab operations: See Greenhab report

Summary of ScienceDome operations:  Currently being cleaned, Washvac assembled, Mission Control organizing Science Dome today

Summary of RAM operations: Not Operational

Summary of health and safety issues:  All Healthy

Questions, concerns and requests to Mission Support:

  1. Request more ATV Fuel. Can probably get through 2 more EVAs before exhausting remaining reserves.
  2. Burnable and Non burnable trash in airlock ready for pickup.

GreenHab Report – December 11th

GreenHab Report

Trisha Randazzo

December 11, 2017

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

·         Outside temperature during work hours:  5 C high -1 C low

·         At 10:20, inside temperature was 40 C. Opened door to cool.

·         Keep door open for most of the day. Closed around 16:00, inside temperature was 20 C.

·         Inside temperature high 15 low 40

·         Inside humidity: 16%


·         Functioning nominally

·         Turns on at 16 C stops at 20 C


·         Not available

Both heating and cooling

·         Nominal

Shade cloth on/off

·         On

Average temperature: (N/A until new sensor is delivered)

·         N/A

Changes to crops:  Note all emergence of seedlings, death of seedlings, etc

·         The original tomato plant is doing well!

·         Two of the four thinned tomato plants are improving!

·         The other two have no change

·         Some of the other tomato plants were starting to wilt, so moved them from the back of the hab to the door where it is cooler

·          One of the beans are starting to produce its first bean!

·         More new seedlings popped up in one of the herb gardens! Probably 20 in  that single planter now!

·         Planted more herb. One herb planter currently has two rows of the following herbs, from left to right (see picture attached to email):

o   Common Sage

o   Oregano

o   Cilantro

o   Parsley

o   Rosemary

o   Dill Weed

Daily water usage for crops:

·         Some of the soil was extra dry

Time(s) of watering for crops:

·         13:15

Morning research observations: (Currently not operational)

·         N/A

Changes to research plants:

·         N/A

Daily watering and amount of water used:

·         2.5 gallons,

Aquaponics:  (Currently not operational)

·         N/A

Narrative:  Any other information you want to share

·         Found the tin of seed packets.

o   Organized the non-empty seed packets alphabetically using cut up cardboard in that tin

o   Created an index of packets which still have seeds

o   Created an empty index of packets to preserve the individual plant care instructions.

·         The current monitor is a fire and carbon monoxide detector. Is there still a desire from mission control to install a separate carbon monoxide detector, perhaps in a different location in the greenhab? See attached pictures for picture of current alarm system.

·         Tested the fire and CO detector, system is nominal.

Support/supplies needed:

·         Mission Control Requests:

o   Request for replanting plan for tomorrow:

§  The planter box closest to the door has about 20-30 seedlings. Permission to replant them into small, individual planters so the box has more space for the rest of the herbs.

Sol Summary – December 11th

Summary Title:  A nominal day!

Mission Status:

Today has been a thankfully nominal day!  Due to our error with comm’s yesterday and not getting our EVA request in on time we stayed at MDRS and concentrated on catching up on HAB maintenance and other activities that have fallen by the wayside in the last few days.

Today we completed our crew training on our ultrasound machine, which enables us to complete our practice runs with the Astronaut trainer developed by Dr. Sczepaniak in the next few days.

We were informed of some likely solar activity by our friendly neighborhood astronomer which led us to perform some highly anticipated solar observation this afternoon.  While we could not detect any prominence we did detect sunspots for the first time.  We are hoping the sun continues in this period of high activity and future solar observation yields some rewarding images.

We continued with our communication experiment and have noticed significant improvement among the crew in terms of numbers of errors and completion times in our building tasks.  We have assembled a lesson’s learned list of communication tips to allow quick and efficient communication among various types of time delay.  We will continue this experiment in various permutations until our mission ends, and then expect to hand over our experiment materials to make this a permanent installation at MDRS that crews can participate in as desired.

We also completed our CPR techniques experiment runs on all crew, analyzing the various ways that more force than body weight can be utilized during a CPR medical emergency.  It was interesting that most crew actually utilized different techniques during this experiment run, with us comparing notes to identify the optimal method.

Our crew has returned to our exercise routine today with all exercise equipment completely assembled.  This has proven to be a crew favorite and we also intend to donate this equipment to MDRS.  We have assembled a stationary exercise bike and a series of medicine balls for weights exercises.

The remainder of the day was filled with HAB maintenance, static water tank transfer, cleaning, some much needed love for our friendly plants in the Greenhab, and assembly of equipment delivered by Mission Control.  We are excited to finish out our mission and eager to get back out into the Martian environment with our 5th Matryoshka experiment site tomorrow.

Thanks for your assistance!

Sol Activity Summary:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. CPR Techniques
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports
  7. EVA Suit Inspection / Maintenance
  8. Washvac Assembly
  9. Power system monitoring
  10. Greenhab Planting
  11. Anesthesia Ultrasound Training
  12. Static Water Tank Transfer

Look Ahead Plan:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. CPR Techniques
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports

Anomalies in work:

1.      HAB Leak from under EVA door continuing (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.  Power System Misconfiguration ongoing.  Generator turned off ~ 10am.  SOC stayed close to 100% all day long.

3.  EVA Suit Malfunctions – See Operations Report

Weather: Cold, clear

Crew Physical Status: All crew in good physical health.


Reports to be filed:



EVA Request

Sol Summary



Support Requested:

1.    ATV Fuel

2.  Burnable / Non Burnable Trash Pickup from Rear Airlock

3.  Permission to replant seedlings into individual planters so herbs have clean soil.  See Greenhab Report for details.

4.  We have an extra CO detector that is marked ‘For Greenhab’.  The current detector in Greenhab is both a Smoke Alarm / CO Sensor.  Please advise whether you would like us to install a redundant CO Sensor in the Greenhab or return it to the HAB and mark it as a replacement part.

3.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken

Journalist Report – December 11th

Crew 184 Journalist Report

Willie Schumann

11 December 2017

Title                            Back on Track

Narrative                  It was our first real weekend on Mars, but there wasn’t much time to lay back and take a rest, because the problems of the past week still had their grip on us. On Friday the generator let us down once again, we shut it down and we prepared for the worst. As we expected another freezing night we started to seal the complete habitat to contain the warmth we have inside. If that still wouldn’t have been enough we considered moving to the science dome, as it is the structure of our Mars mission that is the best insulated. We covered and secured sensitive electrical equipment and send a few thoughts to our loved ones.

We succeeded and had a descent night. I think, the insights we gathered from this emergency event could be useful for the men and women, who will follow us to Mars. But for the time being, we stayed alert, because the problems weren’t resolved. Even if we slept like native Martians, we still had to take care of the generator and the flow of energy at night.

Together with mission control we made the decision to evacuate for the night of Saturday to an emergency pod a few miles south of the Habitat. Equipped with the most relevant life sustaining technologies, perfectly insulated but stripped of the extended possibilities of the habitat this pod served us perfectly for the night, before we could tackle the problem of the generator once another morning provided us with the energy of the sun.

When we returned Sunday morning things had calmed down and we realized, that one source of the problems could have been contaminated coolant in the generator. We decided to flood the coolantInline image 1 tank, to clean the device and refill it with a new mix. Mission control works tirelessly to find additional solutions to the hick-ups in the communication between the battery and the generator and I am sure the earths best scientists are writing already a perfect plan to make our energy infrastructure work not only for us, but for the future crews to inhabit this strange new planet.

Quick flashback to Friday, which was a grand Martian day. We continued our geological Matryoshka project on an EVA consisting of Science Officer Trivedi, Doctor Sczepaniak and me. We went North-East from the habitat and once again the landscape and nature proved to be surprising and new. The location was hidden, but in a mix of satellite imagery, GPS and common sense we finally found Candor Chasma. What a mystical name and a very fruitful location for our scientists.

The surroundings looked very similar to the Sahara Desert on Earth. Deep sandy dunes and mysterious desert plants, complimented by giant stones, that look like they have been dropped from the sky by a higher force. Cunningly some of these rocks have a very large body, but they are only connected to the ground by a small fraction of their bottom parts. Martian rocks seem to defy the logic of physics and there is much to be explored about the origin of these formations.

The crew climbed up to a higher plateau, even when the sun was high and the temperatures in the suits were rising constantly. Trivedi and Sczepaniak really worked as a team and secured a lot of uncontaminated probes. We saved a lot of time compared to former EVA’s and had even time for a film interview at the location in space suits.

After our return to the hab we had much time to care about maintenance., especially because we didn’t have an EVA today. We repaired a few space suits that were a bit out of shape and First Officer Randazzo really dug deep into the soil of the green hab. She had been afraid that after the generator failure, would have effected the temperatures immensely, threatening the livelihood of the young seedlings, she had planted before. But the structure proved its value and all the plants are alive and kicking.

She continued to thin out tomato plants and seed new herbs today to comple-ment the diet plan of the crew. As we were all stuck in the habitat today it becomes apparent, that it is good for the crew members to find refuge in one or another building on the premises to have a little time for themselves. First Officer Randazzo really enjoys the time in the green hab, it is almost like a form of meditation to plant new life into little pots.

Personal Logbook             After a tumultuous few days it was good to come back to the daily routine. There were a few doubts, how everything will continue on Mars for us, but now we can value our time here even more.

After the biggest problems seem to be contained I am looking very forward to go out on the next EVA tomorrow. While the Crew repaired the generator I used the time to shoot some pictures around the habitat. The design of the observatory is a real piece of art and I fell in love with it.

Looking forward to get outside the habitat tomorrow and I will try to capture some scenes, that I haven’t been shot before. I am very anxious what the next days will bring.

We had a great ten days of meals and we are running out of fresh supplies. We have to get creative now for cooking, but I am very confident, that we will create something delicious for the crew, because good food keeps the spirits high.

Thank you very much for your help and attention.

Willie Schumann, Journalist, Crew 184

Copyright © The Mars Society. All rights reserved. | Main Site